OpenStudy (anonymous):

Nancy bought 360 tropical fish for a museum display. She bought 8 times as many goldfish as angelfish. How many of each type of fish did she buy?Which is a system of equations to model the problem if x represents the number of angelfish Nancy bought, and y represents the number of goldfish she bought?

OpenStudy (anonymous):

x – y = 360 y = 8x x + y = 360 y = 8x x + y = 8 y = 360x x + 8y = 360 y = 8x

OpenStudy (anonymous):

Its B.